### v3

@@ 4.2 Results - RQ2: Proposed methodologies for detecting SATD in software development
Automated approaches are characterized by the use of ML, DL, or Transformers and advanced NLP techniques to fully automate the SATD detection process. In order to detect SATD from issue tracking systems, Dai and Kruchten 50 created an automated method that makes use of an ML algorithm, namely of a Naïve Bayes (NB) classifier in addition to other NLP approaches like TF-IDF and TextRank. This method demonstrated the potential for high scalability and integration into software development processes, with promising results in terms of precision, recall, and F1-score. Recently, Li et al. 76 introduced a novel approach to automatically detect SATD from four different sources: source code comments, commit messages, pull requests, and issue tracking systems. The most recent trend in automated SATD detection includes the use of Transformers-based models (e.g., BERT 93, 75, RoBERTa 98, 95, and GPT 99).
Table 9: Textual artifacts Used. Studies that use more than one textual software artifact are highlighted with square brackets
Artifact Paper Reference NLP ML DL TRANSFORMERS
Code comment No. 8, 36, 37, 38, 51, 52, 53, 54, 55, 56, 57, 77, 58, 78, 40, 40, 60, 61, 62, 41, 42, 63, 64, 80, 81, 44, 82, 83, 45, 84, 101, 85, 86, 65, 66, 88, 67, 89, 73, 46, 68, 91, 92, 70, 93, 94, 75, 96, 99, 71, 49, 48, 72, 97,[69, 76, 90, 100] 8, 36, 37, 38, 40, 41, 44, 45, 46, 49, 48 51, 52, 53, 54, 55, 56, 57, 58, 60, 61, 42, 64, 84, 101, 65, 66, 67, 73, 68,[69], 70, 71, 72 77, 78, 79, 62, 63, 80, 81, 82, 83, 86, 88, 89,[76], [90], 91, 92, 94, 96 93, 75, 99,[100], 97
Issue No. 50, 59, 39, 43, 102, 74, 87, 95, [69, 76, 90, 100] 59, 39, 102, 45 50, 84,[69] 74, 87,[76, 90] 95,[100]
Pull request No. 98,[69, 76, 100] - [69] [76] 98,[100]
Commit message No. 47,[69, 76, 100] 47 [69] [76] [100]
Moving forward from the methodological categorization, it is essential to understand the source or type of textual artifacts analyzed in the context of SATD. Table 9 categorizes the various studies on SATD detection based on the types of textual artifacts utilized. Code comments (such as from Java projects, blockchain projects, Dockerfile, and XML files), issues (from issue tracking systems and GitHub issue reports), pull requests, and commit messages contain valuable information for SATD detection, providing varying levels of granularity and relevance depending on the artifact analyzed. The application of NLP, ML, DL, and Transformers-based models across these artifacts shows how methodologies have evolved to handle different levels of complexity and scale.
Code comments are the most commonly analyzed artifact for SATD detection, with NLP, ML, DL, and Transformers models being widely employed (e.g., studies 8, 36, 93). While issues, pull requests, and commit messages are also explored, they are less frequently studied. Early research on SATD primarily relied on NLP to extract patterns from developer language in code comments and issues. However, more recent studies have incorporated ML, DL, and Transformers models (e.g., 76, 93, 98) for a deeper and more automated analysis, especially when dealing with larger and more complex datasets. The increasing adoption of Transformers models reflects the field’s shift towards more sophisticated and scalable SATD detection across various textual artifacts.
Meanwhile, there has been a noticeable shift towards using a more diverse set of textual artifacts for detecting SATD. Furthermore, studies that utilize more than one textual software artifact (e.g., code comments combined with issue trackers or commit messages) are highlighted with square brackets. The increased use of pull requests and commit messages (e.g., 69, 76, 100) as sources of SATD detection in recent studies signals a growing recognition of these artifacts’ value in providing temporal and contextual data on technical debt. As the field moves forward, there is an increasing emphasis on automation and leveraging diverse textual artifacts, which aims to improve the scalability and reliability of SATD detection across varied software projects. The studies that employed multi-artifacts represent an effort to enhance the accuracy and robustness of SATD detection by cross-referencing different sources of developer commentary and interaction.
@@
Answer to RQ2 Over the past decade, there has been a visible shift from manual towards fully automated approaches to improve efficiency and scalability in SATD detection. Additionally, the use of multiple textual artifacts has grown, enhancing the accuracy and robustness of SATD detection by integrating diverse sources from the software development process.
4.3 Results - RQ3: Effectiveness of the proposed approaches regarding their metrics evaluation results
To evaluate the effectiveness of the proposed SATD detection approaches, we examined the performance metrics reported in the studies included in our review. While a subset of the studies did not provide detailed performance results, the majority (49 out of 68) reported on key evaluation metrics such as precision, recall, and F1-score. The following section presents a detailed analysis of these metrics, utilizing boxplots to visualize the performance variability and trends across different SATD detection approaches.
@@ 4.3 Results - RQ3: Effectiveness of the proposed approaches regarding their metrics evaluation results
Out of 68 relevant studies, 17 did not report their performance results, and 2 only provided AUC scores. Therefore, 49 out of the 68 articles included in our review reported performance results using at least one of the common evaluation metrics: precision, recall, or F1-score. We collected all performance results presented in these 49 articles and organized them into a spreadsheet. Rather than focusing solely on the best-performing model from each study, we included the results from all model classifiers. This approach provides a more comprehensive and unbiased evaluation of SATD detection models 103. By considering all models, researchers can capture a broader range of performance variability, which is essential for understanding the factors that influence model effectiveness. Limiting the analysis to only the best-performing models could lead to skewed conclusion, as it overlooks the context in which the models perform and fails to account for suboptimal models that offer important insights into potential limitations or challenges. Moreover, analyzing all models enables a more robust cross-study comparison, allowing for the identification of consistent trends and patterns in model performance across different contexts, datasets, and techniques. This holistic approach strengthens the reliability of the conclusion and ensures their applicability across a wider range of scenarios, rather than being limited to cases of optimal performance.
We use boxplots to demonstrate the effectiveness of various approaches, as they are a useful graphical tool for showing variations between populations without relying on assumptions about data distribution 103. We compiled the effectiveness of various techniques as reported in the original papers, utilizing precision, recall, and F1-score derived from calculations based on the confusion matrix. Through boxplots, we showcase studies on precision, recall, and F1-score in relation to different factors of the models.
@@
  • F1-score = $\frac{2*Precision*Recall}{Precision+Recall}=\frac{2*TP}{2*TP+FP+FN}$
We acknowledge the potential for misinterpretation when aggregating the results of approaches across various studies without considering the diversity of datasets utilized. The rationale behind this methodological choice is to cultivate an inclusive understanding of each model’s performance and versatility across a wide range of contexts and scenarios. This inclusive approach enables the identification of universally effective strategies as well as those that excel under specific conditions, thereby providing a nuanced perspective on the potential applicability, hereby offering a holistic view of their effectiveness in detecting SATD. This methodology aligns with precedents in the field, notably reflected in the systematic literature review, such as the study of Hall et al. 103 and Malhotra 104. This precedent underscores the validity and relevance of cross-dataset comparisons in extracting meaningful insights that transcend the particularities of individual studies. By synthesizing findings across diverse datasets, this review aspires to contribute to the advancement of SATD detection approaches, offering valuable guidance for future research and practical applications in software engineering.
To avoid any misleading interpretations, we divide the process of SATD identification into two distinct tasks: SATD identification and SATD categorization. SATD identification focuses on determining whether a particular entity or instance is present in the text of a software development artifact, classifying it as either SATD or non-SATD. SATD categorization, on the other hand, takes this a step further by not only identifying the presence of SATD but also classifying it into specific types of technical debt. These categories may include labels such as code debt, design debt, requirement debt, and others, based on the linguistic elements found in the text.
@@ 4.3.1 Performance results of SATD identification
The box plot shown in Figure 5 provides a comparative overview of performance across different SATD identification methods, specifically for binary classification, which aims to accurately distinguish between SATD and non-SATD instances. The three key metrics displayed in the box plot—precision, recall, and F1-score—are standard measures for evaluating the effectiveness of classification models. Precision indicates the model’s ability to correctly identify instances as SATD or non-SATD, while recall measures how effectively the model captures all relevant SATD instances. The F1-score, which represents the harmonic mean of precision and recall, offers a balanced evaluation, especially when one metric is prioritized over the other. Together, these metrics provide valuable insights into the performance of SATD detection methodologies in identifying technical debt within software projects.
Figure 5: Boxplot of SATD identification grouped by their approaches
@@
The NLP-based approach demonstrates a relatively high precision of 0.79, indicating that it successfully identifies SATD in a significant number of cases when it classifies an instance as SATD. However, its recall is lower at 0.62, suggesting that it overlooks a considerable portion of SATD instances. The F1-score of 0.66 reflects the imbalance between precision and recall, highlighting that while NLP methods were foundational in early SATD detection, they now face challenges in fully capturing all relevant SATD instances.
@@
In contrast, Machine Learning (ML) models offer a more balanced performance, achieving a precision of 0.73 and recall of 0.74, which results in an F1-score of 0.73. This improvement over NLP methods illustrates ML’s effectiveness in reducing false positives while maintaining good coverage of SATD instances. Historically, ML approaches have represented a significant advancement in SATD detection, effectively bridging the gap between precision and recall.
@@
Deep Learning (DL) models further enhance performance, with precision and recall scores of 0.78 and 0.76, respectively, yielding a balanced F1-score of 0.76. This suggests that DL techniques are better at identifying SATD instances while minimizing false positives. The improvements in both precision and recall demonstrate the advancements deep learning architectures have brought to the complexities of SATD identification, enabling more accurate classification across diverse software contexts.
@@
The Transformers-based approach achieves the best overall performance, with a precision of 0.83—the highest among all methods—and a recall of 0.76, leading to an F1-score of 0.76. These results underscore the superior capability of Transformers models to accurately identify SATD while retrieving a majority of relevant instances. As the latest advancement in SATD detection techniques, Transformers represent the current state-of-the-art, providing both the precision needed to avoid false positives and the recall required to capture a wide range of SATD instances.
@@
Referring to Figure 5, we can further elaborate and illustrate the metric results in more detail, organizing them based on their respective models, as depicted in Figure 6. Each model is evaluated with a specific number of experiments (denoted by “n="), highlighting the range of experiments used to measure their performance.
@@
NLP-based models, particularly Textual Patterns and Annotation Tags, were assessed for their effectiveness. Textual Patterns (n=71) achieved relatively high precision, indicating that this model accurately classifies SATD examples with low false positives. However, its lower recall (0.62) and F1 score (0.66) suggest that it often fails to identify many SATD instances, limiting its overall effectiveness in comprehensive identification. In contrast, Annotation Tags (n=21) performed better in recall and F1 score, indicating a more balanced detection capability, although variability was present, as shown by the wider interquartile range.
@@
ML models exhibited a broader range of performance. Among them, Random Forest (RF) (n=34) and Naive Bayes (NB) (n=37) provided the most consistent results, achieving high precision and F1 scores. This indicates their effectiveness in reducing false positives while capturing a majority of SATD instances. However, models like Support Vector Machine (SVM) (n=13) and k-NN (n=10) showed lower precision and recall, indicating challenges in accurately identifying SATD instances. Notably, Auto-sklearn and LightGBM performed well, though their small sample sizes (n=3 and n=1, respectively) limit the generalizability of their results.
DL models have made significant strides in SATD identification. BiLSTM (n=168), CNN (n=151), and LSTM (n=100) achieved high scores across all metrics, demonstrating consistent precision, recall, and F1 scores. These models effectively handle complex SATD detection tasks while minimizing false positives. However, models like Graph Neural Network (GNN) (n=35), Recurrent Neural Network (RNN) (n=20), and Graph Convolutional Network (GCN) (n=20) exhibited lower performance, particularly in recall, indicating they may miss important SATD instances. Models with smaller samples, such as Ladder Network (LN) (n=10) and Gated Graph Neural Network (GGNN) (n=10), also showed variability in performance, likely due to their limited test sizes.
Transformer-based models represent the most recent advancements in SATD identification. Among these, BERT (n=69), RoBERTa (n=39), and Transformer (n=40) achieved the highest precision, recall, and F1 scores. These models demonstrated remarkable consistency and reliability in identifying SATD instances, with minimal variability, making them leading techniques in SATD detection. Other Transformer models, such as ALBERT (n=18), DeBERTa (n=10), and GPT (n=3), also performed well, though slightly less robustly than the BERT-based models. DistilBERT (n=2) showed promise, but its limited sample size makes it difficult to draw definitive conclusions about its overall performance.
@@
Figure 6: Models performance of SATD identification grouped by each approach
@@ 4.3.2 Performance result SATD categorization
Figure 7 offers insight into the precision, recall, and F1-score of these approaches, focused on various SATD categorization methodologies for multiclass classification, and provides insight into how effectively these compiled approaches differentiate between types like design debt, requirement debt, documentation debt, test debt, defect debt, and others. The metrics provide insights into how effectively these approaches perform in a multiclass classification scenario, where distinguishing between different categories of SATD is more complex than a binary classification.
Figure 7: Boxplot of precision, recall, and F1-score of SATD categorization
@@
NLP approaches demonstrate a low Recall of 0.11 and an F1-score of 0.16, indicating significant difficulty in capturing true SATD instances and missing many relevant examples. This limitation affects their effectiveness in real-world applications, particularly in complex multiclass classification tasks. In contrast, ML approaches show notable improvement, achieving a Precision of 0.75 and a Recall of 0.57, resulting in a balanced F1-score of 0.52. These results suggest that ML models provide better accuracy in SATD classification, although they still miss a considerable number of true positive cases, as indicated by the Recall.
@@
DL approaches perform well across all metrics, with a strong Precision of 0.77 and a high Recall of 0.76. However, the F1-score of 0.62 suggests that while they successfully capture most true positives, there is still potential for improvement in balancing false positives and false negatives. Interestingly, Transformer-based approaches, despite being more advanced, exhibit slightly lower overall performance in this instance, with a Precision of 0.69, Recall of 0.59, and an F1-score of 0.54. While transformers typically excel in complex text classification tasks, these results may reflect the influence of model configuration or data characteristics on their performance in this specific SATD identification task. Nonetheless, transformers still outperform NLP approaches and are comparable to ML models in this context.
@@
Referring to Figure 7, we can provide a comprehensive evaluation of specific types of SATD categorization, organized by their respective models, as shown in Figure 8. Each model is evaluated with a specific number of experiments (denoted by “n="), highlighting the range of experiments used to measure their performance.
@@
Figure 8: Models performance of SATD categorization grouped by each approach
@@
In the NLP group, two approaches, Textual patterns (n=46) and Annotation tags (n=2), were evaluated. The Textual Patterns approach showed highly variable performance, with precision hovering around 0.6, while the recall and F1 scores dropped significantly, indicating that while this method can accurately categorize SATDs in some cases, it struggles to consistently retrieve all SATD instances. The large interquartile range and the presence of outliers highlight the inconsistency of this method in different contexts. In contrast, Annotation tags, although evaluated on only two samples, showed relatively stable and high precision, recall, and F1 scores, indicating strong performance for the multiple cases tested, albeit with limited generalizability. These findings confirm the limitations of NLP techniques in handling the complexities of SATD categorization. While effective in some cases, particularly for high precision, NLP methods often miss critical SATD instances, leading to suboptimal recall and overall lower F1-scores. This aligns with the review’s conclusion that NLP approaches, although foundational in early SATD detection, are no longer sufficient for comprehensive and consistent categorization of SATD.
ML models showed a wide range of results, reflecting the diversity in algorithm complexity and adaptability to SATD categorization. Random Forest (RF) (n=114) and Auto-sklearn (n=108) showed the highest performance, consistently providing high precision and recall, indicating that these models are able to balance false positives and missed SATD examples effectively. However, Naive Bayes (n=17) showed weaker performance, particularly in terms of recall, indicating that it may miss important SATD cases. Other models, such as SVM (n=25), XGBoost (n=50), and Logistic Regression (n=24), showed lower recall and more significant performance variation across the dataset, reflecting their challenges in categorizing SATD consistently.
DL models showed significant improvement over the NLP and ML approaches, specifically in their ability to maintain high precision, recall, and F1 scores across larger data sets. CNN (n=163), CNN-BiLSTM (n=20), and LSTM (n=70) stood out for their consistent and robust performance, with higher recall rates indicating their effectiveness in detecting a wider range of SATD events without significantly increasing false positives. GGNN and Feedforward neural network (FFNN), despite their smaller sample sizes, also showed strong performance, although the small dataset makes it difficult to draw broader conclusion about their generalizability.
Transformers-based models, specifically BERT (n=47), ALBERT (n=24), and RoBERTa (n=13), were the highest performing approaches in SATD categorization. BERT emerged as the best performer, with high precision, recall, and F1 scores across all evaluations, demonstrating its ability to accurately and consistently detect SATDs with minimal false positives and missed events. ALBERT and RoBERTa also performed strongly, although evaluated on smaller sample sizes, their high precision and F1 scores reinforced their value in SATD detection.
@@
From Figures 6 and 8, we can determine the best-performing algorithm for each group based on precision, recall, and F1-score. It is important to note that Figure 6 represents SATD identification aimed at distinguishing SATD from non-SATD (binary classification), while Figure 8 illustrates the results of SATD categorization of specific SATD types (multiclass classification). To select the best-performing algorithm from the boxplot, we primarily focused on the F1-score (dark blue bars) as it offers a balance between precision and recall. We identified the algorithm with the highest median F1-score, indicated by the middle line of the boxplot, and considered its variance (the spread of the box and whiskers), with lower variance suggesting more consistent performance. Additionally, we examined the precision and recall (light and mid-blue bars) to ensure a balance between the two, prioritizing algorithms with high values in both. In addition, we minimized the influence of algorithms with numerous outliers, as they indicate inconsistent performance, even if their median scores were high. Finally, we factored in the n value (number of experiments), giving preference to algorithms tested more frequently, as a higher n suggests more reliable and stable results across different trials.
@@
Table 10: Summary of highest performance results based on SATD detection tasks
Approach Metric Evaluations
Precision Recall F1-score
@@
NLP Annotation tags Annotation tags Annotation tags
ML SMO LR SMO
DL GNN BiLSTM BiLSTM
Transformers GPT Transformer BERT

### v4

@@ 4.2 Results - RQ2: Proposed methodologies for detecting SATD in software development
Automated approaches are characterized by the use of ML, DL, or Transformer and advanced NLP techniques to fully automate SATD detection process. In order to detect SATD from issue tracking systems, Dai and Kruchten 26 created an automated method that uses an ML algorithm, specifically, a Naïve Bayes (NB) classifier, along with other NLP approaches such as TF-IDF and TextRank. This method demonstrated the potential for high scalability and integration into software development processes, with promising results in terms of precision, recall, and F1 score. Recently, Li et al. 30 introduced a novel approach to automatically detect SATD from four different sources: source code comments, commit messages, pull requests, and issue tracking systems. The most recent trend in automated SATD detection includes the use of Transformer-based models (e.g., BERT 101, 84, 31, RoBERTa 107, 104, DistilRoBERTa 109, and ChatGPT 108).
@@
Answer to RQ2 Over the past decade, there has been a visible shift from manual towards fully automated approaches to improve efficiency and scalability in SATD detection.
4.3 Results - RQ3: Synthesis of reported effectiveness metrics for SATD detection techniques
As emphasized in Section 1, a key contribution of this review is the systematic compilation and categorization of reported performance scores across SATD detection techniques. Rather than conducting new experiments or proposing a benchmarking study, we provide a structured cross-technique perspective on how detection approaches have been reported to perform in the literature.
To address RQ3, we synthesized the performance metrics as reported in the selected studies. It is important to note that the synthesized results are limited to within-project evaluations, typically performed using k-fold cross-validation or random splits on the same dataset or project. While some primary studies also reported cross-project or out-of-distribution evaluations (e.g., training on one project and testing on unseen projects), these were deliberately excluded to ensure comparability across studies. Consequently, the reported precision, recall, and F1 scores should be interpreted as reflecting performance under controlled within-project settings, which often yield higher values than cross-project evaluations and may overestimate real-world generalizability.
Following the approach of prior prominent SLRs 112, 113, 114, we synthesized evidence from diverse data sources and models by compiling results reported in the primary studies, rather than conducting new experiments or direct comparisons. Because these studies vary in dataset characteristics, size, and benchmarking protocols, the aggregated results should be interpreted as indicative trends rather than definitive rankings. Where possible, we report the range and central tendency of precision, recall, and F1 scores to illustrate how different approaches have performed in their original contexts.
@@ 4.3 Results - RQ3: Synthesis of reported effectiveness metrics for SATD detection techniques
Out of 74 relevant studies, 17 did not report their performance results, and 3 only provided AUC scores. Therefore, 54 out of the 74 articles included in our review reported performance results using at least one of the common evaluation metrics: precision, recall, or F1 score. We compiled all performance results reported in these 54 articles and organized them in a spreadsheet. Rather than focusing solely on the best-performing model from each study, we included the results from all model classifiers. This approach provides a more comprehensive and unbiased summary of SATD detection models. By considering all models, researchers can capture a broader range of performance variability, which is essential for understanding the factors that influence model effectiveness 112. Limiting the analysis to only the best-performing models could lead to a skewed conclusion, as it overlooks the context in which the models perform and fails to account for suboptimal models that offer important insights into potential limitations or challenges. Moreover, analyzing all models enables a more robust cross-study comparison, allowing for the identification of consistent trends and patterns in model performance across different contexts, datasets, and techniques. This holistic approach strengthens the reliability of the conclusion and ensures its applicability across a broader range of scenarios, rather than being limited to cases of optimal performance.
We used boxplots to visualize the synthesized results because they are useful for showing variation across populations without assuming specific data distributions 112. We extracted the reported effectiveness of each technique from the primary studies, focusing on precision, recall, and F1 scores computed from the confusion matrix. The boxplots present the distributions of these metrics across studies and enable comparisons by model-related factors.
@@
  • F1 score = $\frac{2*Precision*Recall}{Precision+Recall}=\frac{2*TP}{2*TP+FP+FN}$
We acknowledge the potential for misinterpretation when aggregating the results of techniques across various studies without considering the diversity of datasets used. We chose this approach to provide a more comprehensive understanding of each model’s performance and versatility across a wide range of contexts and scenarios. This inclusive approach enables the identification of universally effective strategies as well as those that excel under specific conditions, thereby providing a nuanced perspective on the potential applicability, hereby offering a holistic view of their effectiveness in detecting SATD. This methodology aligns with precedents in the field, as reflected in systematic literature reviews such as those by Hall et al. 112 and Malhotra 113. This precedent underscores the validity and relevance of cross-dataset comparisons in extracting meaningful insights that transcend the particularities of individual studies. By synthesizing findings across diverse datasets, this review aims to advance SATD detection techniques and provide valuable guidance for future research and practical applications in software engineering.
To avoid any misleading interpretations, we divide the process of SATD detection into two distinct tasks: SATD identification and SATD categorization. SATD identification aims to determine whether a particular entity or instance is present in the text of a software development artifact and to classify it as SATD or non-SATD. SATD categorization, on the other hand, takes this a step further by not only identifying the presence of SATD but also classifying it into specific types of technical debt. These categories may include labels such as design debt, architecture debt, and requirement debt, based on the linguistic elements identified in the text.
@@ 4.3.1 Performance results of SATD identification
The boxplot shown in Figure 5 provides a comparative overview of performance across different SATD identification methods, specifically for binary classification, which aims to accurately distinguish between SATD and non-SATD instances. The three key metrics displayed in the boxplot, namely precision, recall, and F1 score, are standard measures for evaluating the effectiveness of classification models. Precision indicates the model’s ability to correctly identify instances as SATD, whereas recall measures how effectively the model captures all relevant SATD instances. The F1 score, which represents the harmonic mean of precision and recall, offers a balanced evaluation, especially when one metric is prioritized over the other. Together, these metrics provide valuable insights into the performance of SATD detection methodologies in identifying technical debt within software projects. Finally, we also report the number of experiments (n) for each technique, giving greater weight to algorithms evaluated more frequently, since larger n values provide more reliable evidence and more stable performance estimates across trials.
Figure 5: Boxplot of SATD identification grouped by their techniques.
@@
The heuristic-based techniques achieve a relatively high precision of 0.72, indicating their ability to correctly identify SATD when classifying an instance as SATD. However, their recall is lower at 0.63, suggesting that they overlook a considerable portion of SATD instances. The F1 score of 0.58 reflects this imbalance, indicating that although heuristic-based methods were foundational in early SATD detection, they now struggle to capture all relevant SATD instances.
@@
In contrast, ML models exhibit more balanced performance, achieving a precision of 0.73 and a recall of 0.73, yielding an F1 score of 0.60. This improvement over heuristic-based methods illustrates ML’s effectiveness in reducing false positives while maintaining good coverage of SATD instances. Historically, ML techniques have represented a significant advancement in SATD detection, effectively bridging the gap between precision and recall.
@@
DL models further improve performance, achieving a precision of 0.78 and a recall of 0.77. This yields a balanced F1 score of 0.76. These results suggest that DL techniques are more effective at identifying SATD instances while minimizing false positives. Improvements in both precision and recall demonstrate how DL architectures have advanced SATD identification by enabling more accurate classification across diverse software contexts.
@@
Transformer-based techniques achieved the best overall performance with a precision of 0.76, the highest recall of 0.87, and an F1 score of 0.78. These results highlight the superior ability of Transformer models to identify most SATD instances while maintaining competitive precision. As the latest advancement in SATD detection, Transformers strike a strong balance between minimizing false positives and maximizing the detection of relevant SATD cases.
@@
Referring to Figure 5, we can further elaborate and illustrate the metric results in more detail, organizing them based on their respective models, as depicted in Figure 6. Each model is evaluated using a specified number of experiments (denoted by “n=”), thereby highlighting the range of experiments used to measure its performance.
@@
Heuristic-based techniques, particularly Textual patterns and Annotation tags, were assessed for their effectiveness. Textual patterns (n=94) achieved relatively high precision, indicating that this model accurately classifies SATD examples with low false positives. However, its lower recall and F1 score suggest that it often fails to identify many SATD instances, limiting its overall effectiveness in comprehensive identification. In contrast, Annotation Tags (n=21) performed better in recall and F1 score, indicating a more balanced detection capability, although variability was present, as shown by the wider interquartile range.
@@
ML models exhibited a broader range of performance. Among them, Random Forest (RF) (n=23) provided the most consistent results, achieving high precision and F1 scores with relatively narrow variability. Logistic Regression (LR) (n=14), while reaching high median values, showed substantial variability across experiments, suggesting less stable performance. Maximum Entropy (ME) (n=46) demonstrated moderately consistent outcomes, though with slightly lower scores compared to RF. In contrast, models such as Support Vector Machine (SVM) (n=48) and k-NN (n=13) exhibited lower precision and recall, indicating difficulty in accurately identifying SATD instances. Naïve Bayes (NB) (n=1) and Sequential Minimal Optimization (SMO) (n=3) were reported only in a few cases, limiting the generalizability of their results. Notably, Auto-sklearn (n=27), LightGBM (n=8), and XGBoost (n=9) also achieved strong results, although their relatively small sample sizes make it difficult to draw robust conclusions.
DL models show heterogeneous performance in SATD identification. BiLSTM (n=170) consistently delivers strong medians, with recall and F1 generally higher than precision and with moderate variability. CNN (n=212) achieves comparable median scores but exhibits the largest dispersion and many low outliers, indicating less stable results across studies. LSTM (n=20) attains solid medians with noticeable spread, while GNN (n=21) also reaches competitive median values but with wide variability and several low-score cases. In contrast, RNN (n=100) trends lower overall, and GCN (n=10) is the weakest, particularly in precision and F1. Among the stronger DL variants, LN (n=79) and GGNN (n=20) stand out with high median performance; GGNN is especially competitive while remaining reasonably consistent despite its smaller sample size.
Transformer-based models represent the most recent advancements in SATD identification. BERT (n=39) and Transformer (n=102) achieved the strongest and most consistent results, with high precision, recall, and F1 scores. ALBERT (n=69) achieves generally strong precision and recall, but its F1 scores vary widely, with several low outliers, making its results less consistent than BERT and Transformer. RoBERTa (n=15) performed well, though with fewer experiments, while DeBERTa (n=18) showed more variability and somewhat lower scores. ChatGPT (n=2) was evaluated in only a very limited number of cases, yielding inconsistent outcomes. DistilBERT (n=5) and Flan-T5 (n=3) achieved promising results but cannot be generalized due to their small sample sizes. CodeBERT (n=44) and DistilRoBERTa (n=62) demonstrated consistently strong performance, with DistilRoBERTa in particular matching the stability of BERT-based models.
@@
Figure 6: Model performance of SATD identification grouped by each technique.
@@ 4.3.2 Performance result SATD categorization
Figure 7 offers insight into the precision, recall, and F1 score of these techniques, focused on various SATD categorization methodologies for multiclass classification, and provides insight into how effectively these compiled techniques differentiate between types like design debt, requirement debt, documentation debt, test debt, defect debt, and others. The metrics provide insights into how effectively these techniques perform in a multiclass classification scenario, in which distinguishing among different categories of SATD is more complex than in binary classification.
Figure 7: Boxplot of precision, recall, and F1 score of SATD categorization.
@@
Heuristic-based techniques demonstrate a relatively high precision of 0.59 but a very low recall of 0.11 and an F1 score of 0.16, indicating significant difficulty in capturing true SATD instances and missing many relevant examples. This limitation affects their effectiveness in real-world applications, particularly in complex multiclass classification tasks. In contrast, ML techniques show notable improvement, achieving a precision of 0.68 and a recall of 0.56, resulting in a more balanced F1 score of 0.52. These results suggest that ML models achieve higher accuracy in SATD classification than heuristics, although they still miss a considerable number of true positives, as indicated by the recall.
@@
DL techniques perform well, achieving a precision of 0.70 and a recall of 0.75, yielding an F1 score of 0.59. This indicates that DL models capture most true positives (high recall) while maintaining reasonable precision; however, the overall balance (F1) remains moderate-higher than ML (0.52) but slightly below Transformer (0.61). Transformer-based techniques, despite being more advanced, achieved a precision of 0.66, a recall of 0.65, and an F1 score of 0.61. While their performance is not markedly higher than that of DL in this setting, Transformers demonstrate consistent, balanced results, outperforming heuristic approaches and ML models and providing comparable effectiveness to DL, with a slightly higher F1 score. While Transformer typically excels in complex text classification tasks, these results may reflect the influence of model configuration or data characteristics on its performance in this specific SATD identification task. Nonetheless, Transformer clearly outperforms Heuristic-based techniques, achieves higher overall scores than ML models, and performs on par with DL approaches, with a slightly better F1 score.
@@
Referring to Figure 7, we can provide a comprehensive evaluation of specific types of SATD categorization, organized by their respective models, as shown in Figure 8. Each model is evaluated using a specified number of experiments (denoted by “n=”), thereby highlighting the range of experiments used to measure its performance.
@@
Figure 8: Model performance of SATD categorization grouped by each technique.
@@
Heuristic-based techniques were evaluated using Textual patterns (n=46) and Annotation tags (n=6). Textual patterns showed highly variable performance, with precision around 0.6 but much lower recall and F1 scores, indicating that while this method can sometimes classify SATDs accurately, it struggles to consistently retrieve all instances. The wide interquartile range and presence of outliers highlight its inconsistency across contexts. In contrast, Annotation tags, although tested in only six cases, showed relatively stable performance, with higher precision, recall, and F1 scores, suggesting stronger performance but limited generalizability. These findings confirm the limitations of heuristic approaches: while sometimes effective for achieving high precision, they often miss critical SATD instances, leading to suboptimal recall and lower overall F1 scores.
ML models exhibited a broad spectrum of results, reflecting their algorithmic diversity. Random Forest (RF) (n=114) and Auto-sklearn (n=114) achieved the strongest performance, with consistently high precision and recall, effectively balancing false positives and missed SATD cases. In contrast, Naïve Bayes (n=17) performed more weakly, especially in recall. Other models, such as Support Vector Machine (SVM) (n=25), Logistic Regression (LR) (n=24), and XGBoost (n=58), showed lower recall and greater variability, indicating challenges in reliably categorizing SATD. Maximum Entropy (ME) (n=18) demonstrated moderate performance, whereas Sequential Minimal Optimization (SMO) (n=8) was tested in only a limited number of cases, thereby limiting its generalizability.
DL models demonstrated clear improvements over heuristic and ML techniques. CNN (n=205) achieved strong and consistent results across metrics, while LSTM (n=70) also performed well but with greater variability. Gated Graph Neural Network (GGNN) (n=20) showed promising outcomes, though based on fewer samples. Feedforward Neural Network (FFNN) (n=4) displayed relatively strong performance, but the extremely small sample size limits conclusions about its broader applicability.
Transformer-based models emerged as the most advanced group. BERT (n=28) achieved consistently high precision, recall, and F1 scores, standing out as the top performer. ALBERT (n=23) also performed strongly, though with more variability, while RoBERTa (n=13) maintained high scores despite its smaller sample size. Additional models, including CodeBERT (n=43) and Flan-T5 (n=22), also demonstrated solid and balanced performance, reinforcing their reliability for SATD detection. DistilRoBERTa (n=2), however, was evaluated on very limited data, making its results less conclusive.
@@
From Figures 6 and 8, we can determine the best-performing algorithm for each group based on precision, recall, and F1 score. It is important to note that Figure 6 represents SATD identification aimed at distinguishing SATD from non-SATD (binary classification), while Figure 8 illustrates the results of SATD categorization of specific SATD types (multiclass classification). To select the best-performing algorithm from the boxplot, we primarily focused on the F1 score (dark blue bars) as it offers a balance between precision and recall. We identified the algorithm with the highest median F1 score, indicated by the middle line of the boxplot, and considered its variance (the spread of the box and whiskers), with lower variance indicating more consistent performance. Additionally, we examined precision and recall (light and mid-blue bars) to ensure a balanced trade-off between the two, prioritizing algorithms with high values for both. In addition, we minimized the influence of algorithms with numerous outliers, as such algorithms indicate inconsistent performance, even when their median scores are high.
@@
Table 10: Summary of highest performance results based on SATD detection tasks.
Technique Metric Evaluations
Precision Recall F1 score
@@
Heuristic Annotation tags Annotation tags Annotation tags
ML RF RF RF
DL CNN BiLSTM BiLSTM
Transformer DistilRoBERTa DistilRoBERTa DistilRoBERTa